OpenStudy (anonymous):

Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

SOMEONE.

OpenStudy (anonymous):

sorry ^-^!!!

OpenStudy (anonymous):

I think u got it though.

OpenStudy (anonymous):

uhh.. you erased it.. I dont know what it is. haha

OpenStudy (anonymous):

Oh well. cya!

OpenStudy (phi):

multiply the bottom equation by -1 that means multiply -1 times each term on both sides can you do that ?

OpenStudy (anonymous):

is this for hw or a quiz/test?

OpenStudy (anonymous):

this is homework.

OpenStudy (anonymous):

but @phi why -1? where am i getting that?

OpenStudy (anonymous):

(3x+2)(x+4) don't listen to him he's wrong.

OpenStudy (anonymous):

foil --> 3x^2+12x+2x+8 --> 3x^2+14x+8

OpenStudy (phi):

you will see in a minute. we are going to make the x term go away

OpenStudy (anonymous):

see, it's the original answer.

OpenStudy (anonymous):

this is a method I've never seen of before ^^^^^^ from phi

OpenStudy (phi):

@daren13354 please do not spam

OpenStudy (phi):

multiply the bottom equation by -1 that means multiply -1 times each term on both sides can you do that ?

OpenStudy (phi):

then add the 2 equations together.

OpenStudy (anonymous):

no this is for substitution. not elimination.

OpenStudy (phi):

ok. (but in the long run, elimination seems easier). but to do substitution on these x − 2y = 14 x + 3y = 9 pick one of the variables to isolate. in this case x looks easier. if we use the first equation: write +2y on both sides of the equal sign and then simplify

OpenStudy (anonymous):

ok so i have x=14+2y

OpenStudy (phi):

now in the 2nd equation x+3y= 9 replace the x with its equivalent value 14+2y simplify

OpenStudy (anonymous):

where do i start first? Combine like terms? subtract 14?

OpenStudy (phi):

what do you have for an equation?

OpenStudy (anonymous):

14+2y+3y=9

OpenStudy (phi):

ok. order does not matter when simplifying. But I would combine the y terms, and also add -14 to both sides of the equation.

OpenStudy (anonymous):

i got y= -1

OpenStudy (phi):

now pick one of the original equations, and replace "y" with -1. simplify to find what x is

OpenStudy (phi):

or use x=14+2y which you found. already to plug in y= -1 to find x

OpenStudy (anonymous):

so... orignal: x-2y=14 now: x-2-1=14

OpenStudy (anonymous):

??

OpenStudy (phi):

x-2y=14 remember -2y is short hand (because people are lazy and don't like to type) -2y is short for -2*y (-2 times y)

OpenStudy (anonymous):

wait...how do i start to solve for x

OpenStudy (phi):

you have x- 2*y= 14 replace y with -1 what do you get?

OpenStudy (anonymous):

like i said up there. x-2y(-1)=14

OpenStudy (anonymous):

correct?

OpenStudy (phi):

replace y. so where you see y, write -1 in x - 2*y=14

OpenStudy (anonymous):

ok now what

OpenStudy (phi):

what do you have for an equation?

OpenStudy (anonymous):

just what do i do now. i have -1 in for y. whats next

OpenStudy (phi):

multiply the 2 numbers. then solve for x

OpenStudy (anonymous):

i got x=16

OpenStudy (anonymous):

........

OpenStudy (phi):

write down the equation you started with, and let's see where you went wrong.

OpenStudy (anonymous):

x-2(-1)=14

OpenStudy (phi):

I am thinking when you started with x - 2* -1 =14 you should think of this as x + -2*-1=14 so when you multiply -2*-1 you get +2 or you could do x - 2*(-1)=14 and 2*-1 is -2: x - -2 = 14 and minus a minus is plus: x + 2=14 does that make sense?

OpenStudy (phi):

now add -2 to both sides to get x

OpenStudy (phi):

or, remember up top where you found "ok so i have x=14+2y" you could replace y with -1 in this equation: x= 14+ 2*-1 2*-1 is -2 so you get x= 14-2 x=12 maybe that is the easiest way.